# MetaSlot: Break Through the Fixed Number of Slots in Object-Centric Learning

### v1

@@ Abstract
Learning object-level, structured representations is widely regarded as a key to better generalization in vision and underpins the design of next-generation Pre-trained Vision Models (PVMs). Mainstream Object-Centric Learning (OCL) methods adopt Slot Attention or its variants to iteratively aggregate objects’ super-pixels into a fixed set of query feature vectors, termed slots. However, their reliance on a static slot count leads to an object being represented as multiple parts when the number of objects varies. We introduce MetaSlot, a plug-and-play Slot Attention variant that adapts to variable object counts. MetaSlot (i) maintains a codebook that holds prototypes of objects in a dataset by vector-quantizing the resulting slot representations; (ii) removes duplicate slots from the traditionally aggregated slots by quantizing them with the codebook; and (iii) injects progressively weaker noise into the Slot Attention iterations to accelerate and stabilize the aggregation. MetaSlot is a general Slot Attention variant that can be seamlessly integrated into existing OCL architectures. Across multiple public datasets and tasks–including object discovery and recognition–models equipped with MetaSlot achieve significant performance gains and markedly interpretable slot representations, compared with existing Slot Attention variants. The code is available at https://anonymous.4open.science/r/MetaSlot.
footnotetext: Correspondence to <hongjia.liu@aalto.fi> and <joni.pajarinen@aalto.fi>.
@@ 2 Method
In this section, we present MetaSlot with (i) First-stage Aggregation for Prototype-guided Pruning; (ii) Second-stage Aggregation with Mask-guided Refinement; and (iii) Prototype update via mini-batch K-means. Notably, the two aggregation modules share weights and are jointly trained. In addition, we include pseudocode in the appendix to provide additional implementation details.
@@ 2.1 Background
2.2 First-Stage Aggregation for Prototype-guided Pruning
@@
Let =$\mathcal{E}=\{\bm{e}_{k}\in\mathbb{R}^{D}\}_{k=1}^{K}$ denote a global codebook of $K$ prototypes. For each intermediate slot vector $\bm{s}_{i}^{\mathrm{mid}}$, we find the index of its nearest prototype and replace it accordingly:
[…]
Duplicate-removal mask.
@@ 2.3 Second-Stage Aggregation with Mask-guided Refinement
In the second-stage aggregation, we refine the pruned prototypes via masked attention with annealed noise injection, producing the final semantically coherent slot set $\bm{S}^{\mathrm{final}}$.
We re-initialize the slot states with the pruned prototype set $\hat{\bm{S}}_{\mathrm{mask}}$ and perform $T$ iterations of attention-based updates. At each step, the raw attention logits are masked by the binary slot mask $smask$, so that only retained prototypes participate in computing the attention-weighted slot increments. To alleviate the “cold-start” misalignment between prior prototypes and current inputs, we also inject progressively isotropic Gaussian noise into the features before each iteration, with variance $\alpha_{t}^{2}$ linearly annealed from $\sigma_{\mathrm{noise}}^{2}$ to zero. This implicit simulated annealing encourages early exploration and late-stage convergence, yielding the final semantically coherent slot set $\bm{S}^{\mathrm{final}}$.
@@
To ensure that only surviving slots steer the refinement, we introduce Masked Slot Attention (MSA). At each iteration, a binary mask $smask$ zeros out rows corresponding to duplicate slots, so the attention-weighted update is computed solely from the retained, semantically meaningful prototype slots. This prevents any duplicate-induced interference. It is worth noting that the MSA shares the same set of weights with the SA used in the first-stage aggregation. For each iteration $t=0,\dots,T-1$ we compute the MSA:
@@
Gradient Truncation and Bi-level Optimization.
Because the vector-quantization (VQ) mechanism truncates gradients–producing instability between Two-stage Slot Aggregation iterations, we stop the gradient flow at the first Slot-aggregation stage. In addition, inspired by the bi-level optimization strategy [29], we further detach gradients during the first $T-1$ iterations of the second Slot-aggregation stage.
@@
2.4 Prototype update via mini-batch K-means.
@@
Prototypes that remain unselected for a predefined timeout window are marked as dead. For each such code we sample a replacement vector $\tilde{\bm{c}}$ from the current mini-batch by choosing the slot that is least similar (cosine distance) to all active prototypes, and reset the dead code via
@@ 3 Related Work
On the theoretical front, a number of studies have offered formal interpretations of object-centric learning (OCL) [54, 27, 55, 56, 57, 58]. In terms of evaluation methodologies, recent works have examined the generalization ability of object-centric representations across various downstream tasks, including visual question answering (VQA) [59], world modeling [24, 25, 60, 61], and video generation [59, 62]. Furthermore, several studies have assessed the robustness of object-centric models under out-of-distribution (OOD) conditions [63, 64, 47].
@@ 4.1 Evaluate on Object Discovery
We integrate MetaSlot into object-centric learning (OCL) frameworks and systematically benchmark it against a range of classic models (Table 1) as well as state-of-the-art models (Table 2) to highlight its performance gains. Concretely, SLATE [22] employs a Transformer decoder for autoregressive reconstruction. DINOSAUR [46] uses an MLP-based hybrid decoder to reconstruct directly in DINO feature space, while VideoSAUR [70] adapts this design to video. SlotDiffusion [42] performs decoding with a conditional diffusion model, and SPOT [71] combines nine permutation-based Transformer decoders with a self-training strategy. Evaluating MetaSlot within each of these heterogeneous decoding paradigms enables a comprehensive assessment of its versatility. We exclude IODINE [49], ISA [72], SAVi [21], SAVi++ [37], and MoTok [73] due to outdated performance or reliance on multi-modal priors that hinder fair comparison under our unified setting
@@
Table 1: Object discovery performance with DINOv2 ViT (s/14) for OCL encoding. The input resolution is 256$\times$256 (224$\times$224). Tfd, MLP and Dfz are Transformer, MLP, and Diffusion [75] for OCL decoding respectively.
@@
SLATE 17.4$\pm$2.9 87.4$\pm$1.7 44.5$\pm$2.2 43.3$\pm$2.4 17.5$\pm$0.6 28.8$\pm$0.3 26.8$\pm$0.3 25.4$\pm$0.3 18.6$\pm$0.1 26.2$\pm$0.8 37.2$\pm$0.5 36.1$\pm$0.4
MetaSlotTfd 40.9$\pm$1.7 92.4$\pm$0.7 49.3$\pm$0.8 48.8$\pm$1.6 18.6$\pm$1.2 33.5$\pm$1.0 28.2$\pm$0.7 26.7$\pm$0.6 20.4$\pm$0.5 30.7$\pm$0.8 39.0$\pm$0.3 37.8$\pm$0.5
DINOSAUR 50.7$\pm$24.1 89.4$\pm$0.3 53.3$\pm$5.0 52.8$\pm$5.2 18.2$\pm$1.0 35.0$\pm$1.2 28.3$\pm$0.5 26.9$\pm$0.5 21.5$\pm$0.7 35.2$\pm$1.3 40.6$\pm$0.6 39.7$\pm$0.6
MetaSlotMlp 64.6$\pm$0.3 89.6$\pm$0.4 55.2$\pm$0.5 54.5$\pm$0.6 22.4$\pm$0.3 40.3$\pm$0.5 29.5$\pm$0.2 27.9$\pm$0.2 32.2$\pm$1.2 43.3$\pm$0.7 43.9$\pm$0.3 42.1$\pm$0.2
SlotDiffusion 66.1$\pm$1.3 82.7$\pm$1.6 54.3$\pm$0.5 53.4$\pm$0.8 17.7$\pm$0.5 28.7$\pm$0.1 27.0$\pm$0.4 25.6$\pm$0.4 17.0$\pm$1.2 21.7$\pm$1.8 35.2$\pm$0.9 34.0$\pm$1.0
MetaSlotDfz 81.9$\pm$0.2 77.6$\pm$1.1 64.2$\pm$0.6 62.8$\pm$0.7 17.7$\pm$0.2 32.2$\pm$0.7 27.2$\pm$0.2 25.8$\pm$0.1 19.0$\pm$0.3 24.1$\pm$0.3 36.5$\pm$0.1 35.3$\pm$0.1
Table 2: Comparison with SOTA methods: SPOT on MS COCO 2017 (images) and VideoSAUR on YTVIS-HQ (videos). All models use a DINOv2 ViT(s/14) backbone. The input resolution is 256$\times$256 (224$\times$224).
@@ 4.3 Interpretability Analysis
As shown in Fig. 3, MetaSlot’s prototype-based initialization yields slots with pronounced semantic consistency and strong object binding—e.g., all slots from prototype 268 correspond to "keyboard" objects, while those from prototype 240 correspond to "umbrella" objects. In contrast, BO-QSA, limited by its fixed number of slots, struggles to achieve such fine-grained prototype binding. Additional results on the VOC dataset are reported in the appendix. Quantitative comparisons under the DINOSAUR [46] decoding framework (Table 4) further confirm that MetaSlot surpasses both SA and BO-QSA on all object discovery metrics, highlighting its superiority in producing disentangled and interpretable slot representations.
@@ 4.4 Ablations
To assess the effectiveness of the key architectural components of MetaSlot, we perform a comprehensive ablation study on the MS COCO 2017 dataset. All experiments adopt the DINOSAUR framework with a DINOv2 ViT (s/14) encoder and fix the number of slots to seven. To validate the contribution of individual design choices, we further evaluate two ablated variants: ’MetaSlot w/o noise’ omits progressively attenuated noise during slot updates. As shown in Fig.5, injecting noise leads to a lower Adjusted Rand Index (ARI, left) and a higher mean boundary overlap (MBO, right), implying faster and more stable slot aggregation. ’MetaSlot w/o mask’ disables the prototype-based masking strategy. Table 5 summarizes performance across three metrics–Foreground ARI (FG-ARI), mean Boundary Overlap (mBO), and mean Intersection-over-Union (mIoU), indicating that each component is pivotal for precise slot-to-object alignment and effective spatial disentanglement.
Figure 4: Ablation study on architectural components. Backbone: DINOv2 ViT (s/14). COCO #slot=7 FG-ARI mBO mIoU MetaSlot w/o noise 39.4$\pm$0.3 28.9$\pm$0.4 27.4$\pm$0.4 MetaSlot w/o mask 38.2$\pm$0.2 28.5$\pm$0.1 26.9$\pm$0.3 MetaSlot 40.3$\pm$0.5 29.5$\pm$0.2 27.9$\pm$0.2        Figure 5: Training curves for MetaSlotMlp with/without progressively attenuated noise. [Uncaptioned image]
@@ Appendix A Method
Input: input features $input$, learnable queries $init$, number of iterations $T$
@@ 15 $\hat{S}_{mask},mask\leftarrow\mathrm{Prune}(\hat{S},idx)$;
16 # Second-Stage Aggregation :
17[0.5em] $S^{final}\leftarrow\hat{S}_{mask}$;
18 for $t=1$ to $T-1$ do
@@ 24 $S^{final}\leftarrow\mathrm{MSA}(S^{final},\bm{i}nput_{noisy},mask)$;
25 # Prototype update :
26[0.5em] $\mathrm{Update}(\mathrm{SG}(S^{final}),mask)$;
@@
Appendix B Limitations
We summarize several limitations of the MetaSlot module in the current work for future researchers to reference and explore in depth:
Codebook Hyperparameters
In this paper, we uniformly set 512 codebook prototypes for all datasets, without systematically studying how this hyperparameter affects model performance. For example, on the VOC dataset, visualization shows that in low-data regimes roughly half of the slots correspond to object instances that occur very infrequently; due to insufficient samples, the prototypes of these “inactive” slots are difficult to learn and visualize effectively, whereas the “active” slots exhibit clear, visualizable object binding. A more thorough investigation of the relationship between codebook size and specific application scenarios may further improve model performance.
Lack of Translation Equivariance
Our current experiments adopt the absolute positional encoding from the original Slot Attention, causing the module to be non-equivariant to image translations. During training, this can lead the codebook prototypes to learn position-dependent noise features. Future work could draw on mechanisms such as ISA [72] to introduce translation equivariance, helping the codebook prototypes converge toward more consistent, position-agnostic representations.
Prototype Optimization via Compositional Generalization
Existing object prototypes focus mainly on overall shape or semantic distribution and have not yet incorporated finer-grained attribute information. By composing and generalizing at the attribute level to build compound prototypes that encode multiple attribute features simultaneously, one may provide more precise initial object cues for downstream tasks. This direction merits exploration in future object-centric learning frameworks.
Out-of-distribution (OOD) Conditions
We have not thoroughly evaluated MetaSlot’s generalization capability across different datasets, nor optimized the matching between codebook prototype distributions and real-world object distributions. Subsequent work could conduct systematic experiments on multi-domain data, aligning codebook prototypes more closely with real data distributions to enhance the model’s cross-sample and cross-domain transfer performance.
Limited Theoretical Understanding
A comprehensive theoretical understanding of how slot initialization influences object grouping is still lacking. Future research is warranted to investigate more principled approaches for constraining and optimizing prototype slot representations—for example, by regularizing them as Gaussian distributions—to enhance both interpretability and representational diversity.
Appendix C Visualization of prototype slots
As shown in Fig. 6 and Fig. 7, we visualize the refined slots $S^{\mathrm{final}}$ corresponding to the initialization prototype slots on the VOC dataset. The codebook contains 512 object prototypes in total. However, since the VOC-trained model is relatively limited in object diversity and scale, many prototype slots receive few or no refined slots assigned to them, making comprehensive visualization infeasible. Therefore, for practical and technical reasons, we focus on the top 20 most active prototype slots—defined as those associated with the largest number of refined slots in the model trained on the VOC dataset. For each selected prototype slot, we randomly sample six refined slots from its assigned set and visualize their corresponding image patches. As shown in the figure, the results clearly demonstrate that the prototype slots exhibit strong concept binding behavior, with refined slots consistently aligned to semantically coherent object categories. These findings are consistent with our theoretical expectations regarding the semantic consistency induced by prototype-guided initialization.
Figure 6: Visualization of slot representations initialized from different prototype slots on the VOC dataset
Figure 7: Visualization of slot representations initialized from different prototype slots on the VOC dataset
Appendix D Experimental Details

### v2

@@ Abstract
Learning object-level, structured representations is widely regarded as a key to better generalization in vision and underpins the design of next-generation Pre-trained Vision Models (PVMs). Mainstream Object-Centric Learning (OCL) methods adopt Slot Attention or its variants to iteratively aggregate objects’ super-pixels into a fixed set of query feature vectors, termed slots. However, their reliance on a static slot count leads to an object being represented as multiple parts when the number of objects varies. We introduce MetaSlot, a plug-and-play Slot Attention variant that adapts to variable object counts. MetaSlot (i) maintains a codebook that holds prototypes of objects in a dataset by vector-quantizing the resulting slot representations; (ii) removes duplicate slots from the traditionally aggregated slots by quantizing them with the codebook; and (iii) injects progressively weaker noise into the Slot Attention iterations to accelerate and stabilize the aggregation. MetaSlot is a general Slot Attention variant that can be seamlessly integrated into existing OCL architectures. Across multiple public datasets and tasks–including object discovery and recognition–models equipped with MetaSlot achieve significant performance gains and markedly interpretable slot representations, compared with existing Slot Attention variants. The code is available at https://github.com/lhj-lhj/MetaSlot.
@@ 2 Method
In this section, we present MetaSlot with (i) First‑stage Aggregation for Prototype‑guided Pruning; (ii) Second‑stage Aggregation with Mask‑guided Refinement; and (iii) Prototype update via mini-batch K-means. Notably, the two aggregation modules share weights and are jointly trained. In addition, we include pseudocode in Appendix A to provide additional implementation details.
@@ 2.1 Background
2.2 First‑Stage Aggregation for Prototype‑guided Pruning
@@
Let $\mathcal{E}=\{\bm{e}_{k}\in\mathbb{R}^{D}\}_{k=1}^{K}$ denote a global codebook of $K$ prototypes. For each intermediate slot vector $\bm{s}_{i}^{\mathrm{mid}}$, we find the index of its nearest prototype and replace it accordingly:
$\mathit{idx}_{i}\;=\;\arg\min_{k}\,\bigl\lVert\bm{s}_{i}^{\mathrm{mid}}-\bm{e}_{k}\bigr\rVert_{2},\qquad\hat{\bm{s}}_{i}\;=\;\bm{e}_{\mathit{idx}_{i}},$ (4)
where $\hat{\bm{s}}_{i}$ is the quantised slot vector, set to the prototype $\bm{e}_{\mathit{idx}_{i}}$.
Duplicate‑removal mask.
@@ 2.3 Second-Stage Aggregation with Mask-guided Refinement
In the second‐stage aggregation, we refine the pruned prototypes via masked attention with annealed noise injection, producing the final semantically coherent slot set $\bm{S}^{\mathrm{final}}$.
We re-initialize the slot states with the pruned prototype set $\hat{\bm{S}}_{\mathrm{mask}}$ and perform $T$ iterations of attention-based updates. At each step, the raw attention logits are masked by the binary slot mask smasksmask, so that only retained prototypes participate in computing the attention-weighted slot increments. To alleviate the “cold-start” misalignment between prior prototypes and current inputs, we also inject progressively isotropic Gaussian noise into the features before each iteration, with variance $\alpha_{t}^{2}$ linearly annealed from $\sigma_{\mathrm{noise}}^{2}$ to zero. This implicit simulated annealing encourages early exploration and late-stage convergence, yielding the final semantically coherent slot set $\bm{S}^{\mathrm{final}}$.
@@
To ensure that only surviving slots steer the refinement, we introduce Masked Slot Attention (MSA). At each iteration, a binary mask smasksmask zeros out rows corresponding to duplicate slots, so the attention-weighted update is computed solely from the retained, semantically meaningful prototype slots. This prevents any duplicate-induced interference. It is worth noting that the MSA shares the same set of weights with the SA used in the first-stage aggregation. For each iteration $t=0,\dots,T-1$ we compute the MSA:
@@
Gradient Truncation and Bi‑level Optimization.
Because the vector-quantization (VQ) mechanism truncates gradients–producing instability between Two-stage Slot Aggregation iterations, we stop the gradient flow at the first Slot-aggregation stage. In addition, inspired by the bi-level optimization strategy [29], we further detach gradients during the first T1T-1 iterations of the second Slot-aggregation stage.
@@
2.4 Prototype update via mini‑batch K-means.
@@
Prototypes that remain unselected for a predefined timeout window are marked as dead. For each such code we sample a replacement vector $\tilde{\bm{c}}$ from the current mini‑batch by choosing the slot that is least similar (cosine distance) to all active prototypes, and reset the dead code via
@@ 3 Related Work
On the theoretical front, a number of studies have offered formal interpretations of object-centric learning (OCL) [54, 27, 55, 56, 57, 58]. In terms of evaluation methodologies, recent works have examined the generalization ability of object-centric representations across various downstream tasks, including visual question answering (VQA) [59], world modeling [24, 25, 60, 61], and video generation [59, 62]. Furthermore, several studies have investigated the robustness of object-centric models under out-of-distribution (OOD) conditions [63, 64, 47]. Other works have explored the use of vector quantization mechanisms to improve object disentanglement and interpretability. For instance, methods such as [65, 66] learn hierarchical, compositional discrete representations that align with objects and their attributes.
@@ 4.1 Evaluate on Object Discovery
We integrate MetaSlot into object-centric learning (OCL) frameworks and systematically benchmark it against a range of classic models (Table 1) as well as state-of-the-art models (Table 2) to highlight its performance gains. Concretely, SLATE [22] employs a Transformer decoder for autoregressive reconstruction. DINOSAUR [46] uses an MLP-based hybrid decoder to reconstruct directly in DINO feature space, while VideoSAUR [72] adapts this design to video. SlotDiffusion [42] performs decoding with a conditional diffusion model, and SPOT [73] combines nine permutation-based Transformer decoders with a self-training strategy. Evaluating MetaSlot within each of these heterogeneous decoding paradigms enables a comprehensive assessment of its versatility. We exclude IODINE [49], ISA [74], SAVi [21], SAVi++ [37], and MoTok [75] due to outdated performance or reliance on multi-modal priors that hinder fair comparison under our unified setting. Appendix E presents a comparative evaluation of MetaSlot against AdaSlot[53], SlotContrast[76], SysBinder[65], and NLoTM[66] across multiple datasets, further demonstrating the effectiveness of MetaSlot.
@@
Table 1: Object discovery performance with DINOv2 ViT (s/14) for OCL encoding. The input resolution is $\times256$ ($\times224$). Tfd, MLP and Dfz are Transformer, MLP, and Diffusion [78] for OCL decoding respectively.
@@
[…]
Table 2: Comparison with SOTA methods: SPOT on MS COCO 2017 (images) and VideoSAUR on YTVIS-HQ (videos). All models use a DINOv2 ViT(s/14) backbone. The input resolution is $\times256$ ($\times224$).
@@ 4.3 Interpretability Analysis
As shown in Fig. 3, MetaSlot’s prototype-based initialization yields slots with pronounced semantic consistency and strong object binding—e.g., all slots from prototype 268 correspond to "keyboard" objects, while those from prototype 240 correspond to "umbrella" objects. In contrast, BO-QSA, limited by its fixed number of slots, struggles to achieve such fine-grained prototype binding. Additional results on the VOC dataset are reported in Appendix B. Quantitative comparisons under the DINOSAUR [46] decoding framework (Table 4) further confirm that MetaSlot surpasses both SA and BO-QSA on all object discovery metrics, highlighting its superiority in producing disentangled and interpretable slot representations.
@@ 4.4 Ablations
To assess the effectiveness of the key architectural components of MetaSlot, we perform a comprehensive ablation study on the MS COCO 2017 dataset. All experiments adopt the DINOSAUR framework with a DINOv2 ViT (s/14) encoder and fix the number of slots to seven. To validate the contribution of individual design choices, we further evaluate two ablated variants: ’MetaSlot w/o noise’ omits progressively attenuated noise during slot updates. As shown in Fig.5, injecting noise leads to a lower Adjusted Rand Index (ARI, left) and a higher mean best overlap (MBO, right), implying faster and more stable slot aggregation. ’MetaSlot w/o mask’ disables the prototype-based masking strategy. Table 5 summarizes performance across three metrics–Foreground ARI (FG-ARI), mean best overlap (mBO), and mean Intersection-over-Union (mIoU), indicating that each component is pivotal for precise slot-to-object alignment and effective spatial disentanglement. In addition, Appendix E presents an analysis of the codebook prototype size, where we observe that increasing the number of prototypes yields only marginal improvements in performance. We also compare MetaSlot models with varying slot counts, and the empirically optimal number of slots aligns with the long-established consensus within the object-centric learning community. Furthermore, we include additional ablations on architectural components, which provide further evidence for the robustness and effectiveness of our model design.
[…]
@@ 5 Conclusion
6 Limitations and Future Directions
Despite its promising results, MetaSlot still faces several limitations that point to fruitful directions for future research. First, the use of absolute positional encoding inherited from the original Slot Attention makes the model non-equivariant to image translations, potentially causing the codebook prototypes to capture position-dependent noise patterns. Future work could leverage translation-equivariant mechanisms such as ISA [74] to promote the emergence of consistent, position-agnostic representations. Second, current object prototypes primarily encode global shape or semantic information while overlooking finer-grained attribute compositions. Enhancing prototype optimization through compositional generalization—that is, constructing compound prototypes integrating multiple attribute-level features—may yield richer and more discriminative object cues for downstream tasks. Third, owing to MetaSlot’s two-stage and iterative optimization design, the framework inherently contains self-supervisory signals. Exploring ways to identify the semantically complete object slots emerging in the second stage and using them to provide weak supervision for the aggregation process in the first stage could further advance the development of variable-slot object-centric learning. Finally, MetaSlot has yet to be extensively evaluated under out-of-distribution (OOD) conditions. Systematic studies across diverse domains, coupled with efforts to better align the learned codebook prototype distributions with real-world object distributions, could further improve the model’s cross-sample and cross-domain generalization.
Acknowledgments and Disclosure of Funding
We acknowledge the support of Finnish Center for Artificial Intelligence (FCAI), Research Council of Finland flagship program. We thank the Research Council of Finland for funding the projects ADEREHA (grant no. 353198). We also appreciate CSC - IT Center for Science, Finland, for granting access to supercomputers Mahti and Puhti, as well as LUMI, owned by the European High Performance Computing Joint Undertaking (EuroHPC JU) and hosted by CSC Finland in collaboration with the LUMI consortium. Furthermore, we acknowledge the computational resources provided by the Aalto Science-IT project through the Triton cluster.
@@ Appendix A Method
40.3ex
Input: input features inputinput, learnable queries initinit, number of iterations $T$
@@ 17 $\hat{S}_{mask},mask\leftarrow\mathrm{Prune}(\hat{S},idx)$;
18
190.5em # Second-Stage Aggregation :
20[0.5em] $S^{final}\leftarrow\hat{S}_{mask}$;
21 for $t=1$ to T1T-1 do
@@ 27 $S^{final}\leftarrow\mathrm{MSA}(S^{final},\bm{i}nput_{noisy},mask)$;
28
290.5em # Prototype update :
30[0.5em] $\mathrm{Update}(\mathrm{SG}(S^{final}),mask)$;
@@
Appendix B Visualization of prototype slots
Figure 6: Visualize slot representations initialized from different prototype slots on the VOC dataset.
As shown in Fig. 6, we visualize the refined slots $S^{\mathrm{final}}$ corresponding to the initialization prototype slots on the VOC dataset. The codebook contains 512 object prototypes in total. However, since the VOC-trained model is relatively limited in object diversity and scale, many prototype slots receive few or no refined slots assigned to them, making comprehensive visualization infeasible. Therefore, for practical and technical reasons, we focus on the top 20 most active prototype slots—defined as those associated with the largest number of refined slots in the model trained on the VOC dataset. For each selected prototype slot, we randomly sample six refined slots from its assigned set and visualize their corresponding image patches. As shown in the figure, the results clearly demonstrate that the prototype slots exhibit strong concept binding behavior, with refined slots consistently aligned to semantically coherent object categories. These findings are consistent with our theoretical expectations regarding the semantic consistency induced by prototype-guided initialization.
Appendix C Experimental Details